Main Logbooks page LST commissioning Shift checklists Safe state checklists Maintenance activities LST2-4 Commissioning Old logbooks
GENERAL CAMERA DRIVE ENERGY AMC
  LST1 Drive ELOG  ELOG logo
Message ID: 250     Entry time: Thu Apr 18 21:07:20 2019
Author: Drive Team- Ino&Thierry&Alessandro 
Type: Telescope Operations 
Category:  
Subject: TESTS Parkout Parkin for fixing the bolt issue 
Status:  
  • Different ParkOut and Parkin done with different values for testing the torque limitation. 
    • By decreasing the torque and the position , the parkin failed because we are not in the CSS locking sensor (in locking position)
    • By setting back the original values, th parkin is ok but the bolt cannot be properly inserted.
ELOG V3.1.5-fc6679b